OpenStudy (anonymous):

Simplify 14 minus 7 times x all over x minus 2.

OpenStudy (anonymous):

\[\frac{ 14-7x }{ x-2 }\] like this?

OpenStudy (anonymous):

...if so....

OpenStudy (anonymous):

\[\frac{ 14-7x }{ x-2 }=\frac{ 7(2-x) }{ x-2}\] then you cross out 2-x and x-2 to get (-1) so basically you get 7(-1)=-7